a food packet is dropped from a helicopter rising up with vel.4m/s.the vel of the particle after 3s will be

Grade:11

1 Answers

Arun
5 years ago
A =9.8m/s^
u=4m/s
t=3
v = u +at
v= 4+9.8×3
7.26 approx is the answer
h = 7m/s
 
Regards
Arun (askIITians forum expert)